Anorganic compound has wavelength max = 480nm. The absorbance

of it solution was found to be 0.560. If an unknown

solution of this

of this compound gave

an absorbance of 0.750

calculate the concentration​


1
Expert's answer
2021-02-10T02:35:49-0500

"y=mx + b"

Y=0.750

X= concentration

m=480×10-9

b=0.56

mx=0.75- 0.56= 0.19

X=0.19/480×10-9

x= 3.96×105